Liquid octane reacts with gaseous oxygen gas (02) to produce gaseous carbon dioxide (CO2) and gaseous water (H2O). What is thLiquid octane reacts with gaseous oxygen gas to produce gaseous carbon dioxide and gaseous water . What is the theoretical yield of carbon dioxide formed from the reaction of of octane and of oxygen gas? Be sure your answer has the correct number of significant digits in it.
Answer #1

Molar mass of C8H18 = (8 \times 12.01 ) + ( 18 \times 1.0079 ) = 114.22 g / mol

Molar mass of O 2 = ( 2 \times 16.00 ) = 32.00 g / mol

Molar mass of CO 2 = 12.01 + ( 2 \times 16.00 ) = 44.01 g /mol

Firstly, we will find limiting reactant.

Consider a given reaction, 2 C8H18 (l) + 25 O 2 (g) phpyJJEKo.png 16 CO 2 (g) +18 H2O (g)

From reaction,2 mol C8H18 \equiv 25 mol O 2

\therefore 2 \times 114.22 g C8H18\equiv 25 \times 32.00 g O 2

\therefore 36.6 g C8H18 \equiv (25 \times 32.00 \times 36.6 / 2 \times 114.22 ) g O 2

36.6 g  C8H18  \equiv 128.17 g O 2

For complete reaction of 36.6 g octane , 128.17 g oxygen is required. Provided amount of oxygen is 202 g. Hence, oxygen is excess reactant and octane is limiting reactant . Yield of carbon dioxide will depend on amount of octane.

From reaction,2 mol C8H18 \equiv 16 mol CO 2

\therefore 2 \times 114.22 g C8H18 \equiv 16 \times 44.01 g CO 2

\therefore 36.6 g C8H18 \equiv (16 \times 44.01 \times 36.6 / 2 \times 114.22 ) g CO 2

36.6 g C8H18 \equiv 112.8 g CO 2

Answer : Theoretical yield of carbon dioxide = 113 g